# New Online Communities: Graph Deep Learning on Anonymous Voting Networks to Identify Sybils in Polycentric Governance

### v4

@@ 6.2.1 Defining Sybils and Deanonymizing Social Networks
We approach the challenge of identifying sybils as a graph signal problem where we exploit poor or low k-anonymity (wallet re-use) and use distinguishing metadata to construct higher dimensional representations to attack high k-anonymous (“singleton”) sybils. Sweeney’s [50] foundational theory of k-anonymity was initially designed to protect privacy by ensuring that any released data could not be traced back to less than k individuals. In the simplest terms, the theory is about blending in to stay anonymous. Some of our dataset is k=1 (singletons): 71% of unknown users reuse wallet addresses (see Figure 4), which we capture in our graph signal as multi-edge connectivity. These are the easier cases and despite a significant impact on privacy, most users on most chains reuse wallets (e.g., in Bitcoin, wallet reuse has hovered around 50-75%; (see Figure 5 from [7]). As explained below, our method to overcome k-anonymous datasets is deceptively simple: we generate node vector embeddings and cluster the high dimensional embeddings to identify similar, unknown voters. Importantly, since we generate (higher dimensional) node embeddings for all nodes with features, we bypass the protections of k-anonymity, making it possible to spot sybils more effectively.
Figure 4: Histogram of Wallet Reuse in Snapshot Voting Dataset. 71% of voters are $\geq$1.
@@
Our method is far from novel, but it considers new contexts and uses sophisticated deanonymization techniques only recently made possible with advances in machine learning and the availability of powerful computational resources.
Prior research, like Backstrom, Dwork, and Kleinberg [4] tackled social network deanonymization, but (like Sweeney) used a dataset that was anonymized after-the-fact. The advantage of these approaches is that the method is fully supervised and has a clear objective. Unfortunately, such methods are not practical against real datasets that achieve their anonymity by cryptographically deriving new addresses. But their methods fit their goals: they propose two graph-based techniques for deanonymization, an active one that requires the prior construction of a subgraph of attacking nodes (prior collusion) to reveal information about neighbours (by walking the graph or making strategic cuts), and a passive method that relies on graph isomorphisms. Narayanan and Shmatikov’s more recent work on social network deanonymization [36] draws inspiration from Backstrom et al., but used graph isomorphisms with auxiliary graphs to link anonymous identities with public identities, a more information rich context and a method that requires a small number of ‘seed’ nodes present in both graphs.
@@
The most effective research on deanonymizing Bitcoin transactions uses real labelled datasets and intuitive machine learning algorithms (Akcora et al. [1] and Paquet-Clouston et al. [40]) or graph deep learning (Zhou et al. [59]), but these studies still have significant limitations. Paquet-Clouston et al. [40] attempt to trace ransomware payments by constructing graphs from 7222 labelled seeds originating from a convenience sample of known ransomware addresses. They construct an address graph and a cluster graph, the latter uses the logic of sybils to combine nodes that have shared inputs (multiple-input clustering heuristics). However, their analysis is purely descriptive and does not attempt to make predictions. Akcora et al. [1] also investigate ransomware transactions, use convenience sample of three datasets (which includes the Paquet-Clouston et al. dataset), and construct a transaction graph and an address graph. They use multiple heuristics, develop graph features, use DBSCAN and k-means algorithms to cluster addresses and a naive similarity search. Finally, Zhou et al. [59] distinguish between address clustering and account identification, the former we call ‘sybil identification’ and the latter ‘deanonymization.’ They construct a trimmed graph of Ethereum transactions, develop node features, and train a graph neural network (GNN) to search subgraph node embeddings using a hierarchical (weighted) graph attention autoencoder. As is typical, they use several pooling layers and non-linear activation layers (LeakyReLu). They use cross-entropy loss for the classifier. They improve the performance of the default GNN by using contrastive subgraph learning. This self-supervised methodology, while in a different and somewhat easier context, is similar to the approach detailed below. The biggest difference is that they classified nodes directly using the GNN, whereas we use develop the embeddings first and then classify using a embeddings search. The second step is necessary to detect sybils because sybils are by definition unknown and with limited connectivity to other voter nodes. Thus we do not have labels for sybils and we cannot a priori determine the number classes for self-supervised learning. A straightforward classification task will not be able to identify sybils.
@@
In the context of decentralized governance, however, sybils are identified using our more relaxed definition that circumvents k-anonymity. In this context, our method for identifying sybils is a weak form of deanonymization but one that reveals no personally identifiable information and unlike most previous methods does not require auxiliary or labelled graphs (although it can benefit from such information). Supervised techniques for deanonymization will not effectively identify sybils because, in a social graph characterized by strong anonymity, sybils may create a single, disconnected edge and thus lack sufficient connectivity to be identified by walk-based deanonymization approaches. Naive representations of sybils contain only 1 bit of structural information. Indeed, we assume that a majority of sybils are singletons. The existence of a labelled subgraph of known voters, linked to Ethereum Name Service (ENS) accounts, provides little help in identifying isolated sybils, for the same reason.
@@
To overcome this limitation, we define a sybil as a cluster of “similar” nodes. Analytically, this makes sense; if nodes have similar higher dimensional representations and hidden features we can conclude that they are either representative of the same voting person or at least representative of voters acting in concert with one another (perhaps conducting a governance voting attack).
@@ 6.3   Experimental Sybil Identification Methodology
The complete dataset covers governance voting for thousands of DAOs between 2020-07-18 and 2023-02-27 (959 days, 2.6 years), which covers approximately 7.2 million voting transactions. However, naive counts of DAOs, voters, and their governance proposals are ineffective due to noise. For instance, upon manual inspection the dataset reveals many registered DAOs that failed to launch or were clearly set up as tests, and many governance proposals have durations far too short to be practical (see Figure 6).
@@
We found evidence of extensive wallet address re-use, where a single voter makes multiple votes from the same pseudonymous wallet address (in the snapshot.org dataset, approximately 71% are reused wallets). In a traditional financial setting, this is considered poor security behaviour (resulting in low k-anonymity). However, in a polycentric governance context, this is not necessarily evidence of lax security practices but rather the demands of a different social context.
@@
We performed typical Extract, Transform, and Load (ETL) practices on our dataset. After querying the snapshot.org API, we stored the dataset in a relational database and used a SQL command to order voting transactions by date. After sanitizing and checking the data, we constructed the voting graph. On the voting graph, we set node and edge attributes, including the amount of ‘voting power’ used for each vote on a proposal, UNIX timestamps, and persistent IDs linked to Ethereum Name Service IDs, when available. The resulting voting graph constructed from the snapshot.org dataset is available for inspection and statistical analysis (see Github). It is trivial to calculate statistical measurements like density and degree distribution, centrality measures like betweenness and eigenvector centrality, and community detection (there are no communities).
6.3.1 Graph Signal Filter Analysis
@@
For our dataset of 7.2m votes, using the PyGSP Python library, we constructed several wavelet graph signal filters to better understand the frequency components of voter time series.
Exponential Window (Expwin) Filter
@@
The Expwin filter, characterized by an exponential decay, is a low-pass filter with a frequency response expressed as:
@@
Figure 7: Spectral Graph Exponential Window Transform Histogram.
Spectral Graph Wavelet Transform (SGWT)
Spectral Graph Wavelet Transform (SGWT) is used to analyze graph signals by computing the Laplacian operator $L$, which encodes the connectivity structure of the graph, and its eigenvalues and eigenvectors, analogous to the classical Fourier transform. When a signal $x$ on a graph undergoes SGWT using a filter $g_{i}$, it is decomposed into components corresponding to different frequency bands captured by the wavelet filters (see Figure 8). Mathematically, the transform is given by:
@@
Meyer Wavelet Filter Bank
The Meyer Wavelet filter bank provides multi-resolution analysis and is a tight frame, suitable for decomposing signals into different scales. Its general form in the spectral domain is:
@@
The Expwin filter is best for capturing overall smoothness, while the Meyer wavelet strikes a balance between smoothness and localization. The SGWT with various filters can provide a comprehensive understanding of the signal’s structure on the graph. High coefficients in the SGWT indicate regions where the signal strongly resonates with the filter, pointing to significant features at different frequency bands.
6.3.2 ‘Follow the money’ Sybil Chain Subgraph Depth-First Search
Using the strict definition of sybils [14], we can query and walk external graphs to search for sybil chains to support our self-supervised method. A sybil chain is a path across the graph that connects two or more unknown voters (singletons or k-anonymous nodes). To find such paths, we first construct multiple subgraphs. Specifically, we query the Covalent API for each wallet transaction on the origin token’s chain, construct subgraphs for each voter and connect these back to the original voter graph. Then, with this much larger graph we conduct a breadth-first search (BFS) for sybil chains. We use the cuGraph Python library for GPU BFS. We conclude that any path between two or more previously unknown voters is a sybil chain (a multiple-input clustering heuristic).
Figure 10: Examples of external subgraphs before attachment to voting graph.
Variants of this ‘follow the money’ approach are typical in the research, but as we discovered experimentally, without constructing exceptionally large graphs the likelihood of finding sybil chains is very low (even in poor k-anonymous datasets, like Bitcoin, Ethereum, and Snapshot), because isomorphisms between induced subgraphs are rare in anonymous settings. For testing purposes, we include a random fake sybil in the dataset and identify it.

### v6

@@ 5.2.2 Defining Sybils and Deanonymizing Social Networks
We approach the challenge of identifying Sybils as a graph signal problem where we exploit poor or low k-anonymity (wallet re-use) and use distinguishing metadata to construct higher dimensional representations to attack high k-anonymous (“singleton”) Sybils. Sweeney’s [56] foundational theory of k-anonymity was initially designed to protect privacy by ensuring that any released data could not be traced back to less than k individuals. In the simplest terms, the theory is about blending in to stay anonymous. Some of our dataset is k=1 (singletons), however 71% of unknown users reuse wallet addresses (see Figure 6), which we capture in our graph signal as multi-edge connectivity (our target users are ‘unknown’ in the sense of having no Ethereum Name Service registration). These are the easier cases and despite a significant impact on privacy, most users on most chains reuse wallets (e.g., in Bitcoin, wallet reuse has hovered around 50-75%; see Figure 7 from [7]). As explained below, our method to overcome k-anonymous datasets is deceptively simple: we generate node vector embeddings and cluster the high dimensional embeddings to identify similar, unknown voters. Importantly, since we generate (higher dimensional) node embeddings for all nodes with features, we bypass the protections of k-anonymity, making it possible to spot Sybils more effectively.
Figure 6: Histogram of Wallet Reuse in Snapshot Voting Dataset. 71% of voters are k$>$1.
@@
Our method is far from novel, but it considers new contexts and uses sophisticated deanonymization techniques only recently made possible with advances in machine learning and the availability of powerful computational resources. Since our research objective synthesizes techniques and concepts from numerous fields, the methods are grounded in data science and digital experiments.
Prior research, like Backstrom, Dwork, and Kleinberg [4] tackled social network deanonymization, but (like Sweeney) used a dataset that was anonymized after-the-fact. The advantage of these approaches is that the method is fully supervised and has a clear objective. Unfortunately, such methods are not practical against real datasets that achieve their anonymity by cryptographically deriving new addresses. But their methods fit their goals: they propose two graph-based techniques for deanonymization, an active one that requires the prior construction of a subgraph of attacking nodes (prior collusion) to reveal information about neighbours (by walking the graph or making strategic cuts), and a passive method that relies on graph isomorphisms (i.e., ‘partitioning’ rather than ‘clustering’). Narayanan and Shmatikov’s more recent work on social network deanonymization [41] draws inspiration from Backstrom et al., but used graph isomorphisms with auxiliary graphs to link anonymous identities with public identities, a more information rich context and a method that requires a small number of ‘seed’ nodes present in both graphs.
@@
The most effective research on deanonymizing Bitcoin transactions uses real labelled datasets and intuitive machine learning algorithms (Akcora et al. [1] and Paquet-Clouston et al. [45]) or graph deep learning (Zhou et al. [65]), but these studies still have significant limitations. Paquet-Clouston et al. [45] attempt to trace ransomware payments by constructing graphs from 7222 labelled seeds originating from a convenience sample of known ransomware addresses. They construct an address graph and a cluster graph, the latter uses the logic of Sybils to combine nodes that have shared inputs (multiple-input clustering heuristics). However, their analysis is purely descriptive and does not attempt to make predictions. Akcora et al. [1] also investigate ransomware transactions, use convenience sample of three datasets (which includes the Paquet-Clouston et al. dataset), and construct a transaction graph and an address graph. They use multiple heuristics, develop graph features, use DBSCAN and k-means algorithms to cluster addresses and a naive similarity search.
Finally, Zhou et al. [65] attempt Ethereum address clustering, or what we would call ‘Sybil identification.’ They construct a trimmed graph of Ethereum transactions, develop node features, and train a graph neural network (GNN) to search subgraph node embeddings using a hierarchical (weighted) graph attention autoencoder. As is typical, they use several pooling layers and non-linear activation layers (LeakyReLu). They use cross-entropy loss for the classifier. They improve the performance of the default GNN by using contrastive subgraph learning. This self-supervised methodology, while in a different context, is similar to the approach detailed below. The significant difference is that they classified nodes directly using the GNN, whereas we use develop the embeddings first and then classify using a embeddings search, an encoder-decoder model. The second step is necessary to detect Sybils because Sybils are by definition unknown and with limited connectivity to other voter nodes.
@@
In the context of decentralized governance, however, Sybils are identified using our more relaxed definition that circumvents k-anonymity. In this context, our method for identifying Sybils is a weak form of deanonymization but one that reveals no personally identifiable information and unlike most previous methods does not require auxiliary or labelled graphs (although it can benefit from such information). Supervised techniques for deanonymization will not effectively identify Sybils because, in a social graph characterized by strong anonymity, Sybils may create a single, disconnected edge and thus lack sufficient connectivity to be identified by walk-based deanonymization approaches. Naive representations of Sybils contain only 1 bit of structural information. Indeed, we assume that a majority of Sybils are singletons. The existence of a labelled subgraph of known voters, linked to Ethereum Name Service (ENS) accounts, provides little help in identifying isolated Sybils, for the same reason (but can be used to validate clusters).
@@
To overcome this limitation, we define a Sybil as a cluster of ‘similar’ nodes, and singletons are noise. Analytically, this makes sense; if nodes have similar higher dimensional representations and hidden features we can conclude that they are either representative of the same voting person or at least representative of voters acting in concert with one another (perhaps conducting a governance voting attack).
@@ 5.3   Experimental Sybil Identification Methodology
The dataset covers governance voting for roughly 35,000 DAOs between 2020-07-18 and 2023-02-27 (959 days, 2.6 years), which covers approximately 7.2 million voting transactions. However, naive counts of DAOs, voters, and their governance proposals are ineffective due to noise. For instance, upon manual inspection the dataset reveals many registered DAOs that failed to launch or were clearly set up as tests, and many governance proposals have durations far too short to be practical (see Figure 8).
@@
We found evidence of extensive wallet address re-use, where a single voter makes multiple votes from the same pseudonymous wallet address (in our dataset, approximately 71% are reused wallets). In a traditional financial setting, this is considered poor security behaviour (resulting in low k-anonymity). However, in a polycentric governance context, this is not necessarily evidence of lax security practices but rather the demands of a different social context.
@@
We performed typical Extract, Transform, and Load (ETL) practices on our dataset. After querying the snapshot.org API, we stored the dataset in a relational database and used a SQL command to order voting transactions by date. After sanitizing and checking the data, we constructed the voting graph. On the voting graph, we set node and edge attributes, including the amount of ‘voting power’ used for each vote on a proposal, UNIX timestamps, and persistent IDs linked to Ethereum Name Service IDs, when available. The resulting voting graph constructed from the snapshot.org dataset is available for inspection and statistical analysis (see Github).
5.3.1 ‘Follow the Money’ Sybil Chain Subgraph Search
Our dataset captures the “off-chain” voting activities of thousands of DAOs on the governance platform snapshot.org (setting these DAOs apart from more autonomous DAOs, like MakerDAO, that use “on-chain” voting to directly modify organizational parameters). Tracing a token vote across its origin chain would clearly be informative and useful for our deep learning method, but this feature proves challenging to engineer. Recall that off-chain voting exists in large part to address high gas fees associated with main chain transactions (Ethereum or otherwise); thus, off-chain voting leaves no transaction chains to follow. Instead, the process of identifying a “Sybil chain” would require pinpointing the specific token(s) ‘used’ at the time of the vote (used in the sense that they are counted by specific voting ‘strategies’), induce an n-deep path on the origin chain, and conduct a breadth-first search to connect two or more previously unknown voters. However, despite its potential to enhance our self-supervised approach, the problems associated with walking origin graphs are significant, such that implementation would entail addressing a distinct set of issues and a different research agenda.
Thus, naively tracing token transactions is thwarted by robust privacy protections inherent in the various origin chains. To adapt to these constraints and for experimental purposes, we insert a randomly generated fake Sybil in the dataset. This controlled element is then identified as part of our testing protocol. The inclusion of this artificial Sybil serves as a test case, assessing the efficacy of our methods under controlled conditions and providing insights into the potential and limitations of our approach in dealing with real-world scenarios.
However, variants of this ‘follow the money’ approach [8] have been successfully developed, refined, and even commercialized. For instance, although no longer maintained, BlockSci was the benchmark for Bitcoin taint analysis and forensics; today, many commercial platforms offer forensics capabilities (Chainalysis is currently a market leader).
Even in the ‘simplest’ case of Bitcoin, privacy preserving blockchains present a number of obstacles to forensic, especially ‘taint’ analysis. Rather straightforwardly, since Bitcoin uses an unspent transaction output (UTXO) accounting model, it is possible to determine when wallet addresses that require the exact same knowledge to spend can be considered equivalent (BlockSci’s “equivalent addresses”), which is a type of direct Sybil identification. However, even when assuming users are not following privacy-maximizing address rotation practices (about half of Bitcoin transactions), tracing financial flows still requires linking inputs and outputs, addressing “change” wallets, and surmounting coinJoin transactions and coin mixing services.
By combining multiple heuristics with supervised classification tasks, [32] were able to improve link prediction across Bitcoin transactions, besting BlockSci’s static and dynamic heuristics. Indeed, Meiklejohn’s corpus of research on the co-spend (or multiple input) clustering heuristic [37] and more complex “peel chains” [32] exemplify the practical challenges of linking Bitcoin transactions. In another recent work, Möser and Narayanan [40] improve on earlier de-anonymizing techniques by refining existing heuristics (including multiple input and change transactions) and using a Random Forest classifier to cluster similar transactions. Like ours, their method uses self-supervised learning (bootstrapped from a collection of heuristics), but uses a more conventional — lower dimensional — clustering method. In addition to heuristics, they engineer a number of fingerprinting features. They report impressive ROC scores but the method is limited by the numerous dataset exclusions required to prevent cluster collapse (e.g., singletons cannot be identified). Notably, this form of “address clustering” is simply another definition of Sybil identification. And finally, while none of these approaches attempt to dissect coinJoin and mixing techniques, the ‘Boltzmann’ coefficient measures the difficulty of predicting links by computing the entropy of Bitcoin transactions. A similar research challenge remains for Ethereum and other chains.
5.3.2 Graph Time Signal Analysis
@@
For our dataset of 7.2m votes, using the PyGSP Python library, we constructed several wavelet graph signal filters to better understand the frequency components of the voter time series. Wavelet transforms offer a powerful mechanism for decomposing time series data into a spectrum of frequencies across different scales, enabling the extraction of nuanced, multi-scale features that encapsulate both the frequency and temporal information inherent to the signal
@@
First, the Expwin filter, characterized by an exponential decay, is a low-pass filter with a frequency response expressed as:
@@
Second, Spectral Graph Wavelet Transform (SGWT) is used to analyze graph signals by computing the Laplacian operator $L$, which encodes the connectivity structure of the graph, and its eigenvalues and eigenvectors, analogous to the classical Fourier transform. When a signal $x$ on a graph undergoes SGWT using a filter $g_{i}$, it is decomposed into components corresponding to different frequency bands captured by the wavelet filters (see Figure 11). Mathematically, the transform is given by:
@@
And finally, the Meyer Wavelet filter bank provides multi-resolution analysis and is a tight frame, suitable for decomposing signals into different scales. Its general form in the spectral domain is:
@@
The Expwin filter is best for capturing overall smoothness, while the Meyer wavelet strikes a balance between smoothness and localization. The SGWT with filters helps describe the signal’s structure on the graph. High coefficients in the SGWT indicate regions where the signal strongly resonates with the filter, pointing to significant features at different frequency bands, which can be exploited by deep learning.
In the pursuit of elucidating the complex temporal dynamics of DAOs, we engineer and embed spectral time series data, use wavelet transforms with unsupervised clustering techniques to reduce the high dimensional signal and visualize with UMAP. This multi-resolution analysis is particularly useful for highlighting non-linear patterns and dynamics obscured in raw time series data, making wavelets a suitable preprocessing step for uncovering spectral time embeddings.
DRAFT: INSERT DIAGRAMS
Following the wavelet transformation, clustering algorithms are used to group vectors based on similarity of their spectral and temporal features, revealing distinct patterns or behaviors within the dataset. Clustering is important for understanding the inherent structure and relationships within the data, providing a basis for further analysis and visualization. Finally, UMAP is used to project the high-dimensional clusters into a lower-dimensional space, preserving the essential topological structure of the data. This dimensionality reduction enables a visual exploration of the clustered spectral time embeddings and underscores the relationships and distinctions between identified patterns. The amalgamation of wavelet-based feature extraction, clustering, and UMAP visualization presents an experimental framework for analyzing time series data, helping unearth and visualize complex spectral time dynamics that would otherwise remain concealed.
DRAFT: INSERT DIAGRAMS
5.3.3 Wallet Fingerprint Signal Analysis
While we cannot directly validate Sybils by tracing origin chain transactions, we can use historical wallet balances to “fingerprint” voters. Using the snapshot.org API, we constructed a graph of voters for ten DAOs of varying sizes and augmented each voter with unique wallet token balances at the time of the vote. We used the Covalent Unified API to collect token tickers and their associated historical balances and attached these ‘coins’ to each voter. From Figures 14 and 14, we see many unique tokens in voter wallets and observe that voters share wallet token sets, with a long tail of shared, distinct voter sets. If we then dot plot the token sets and colormap their normalized values, we see clear evidence of extractable information for each wallet defined by the uniqueness of wallet token sets (Figure 15).
Figure 13: Histogram of Unique Wallet Tokens.
Figure 14: Histogram of Unique Wallet Token Sets.
Figure 15: Dot Plot of Voter Wallet Token Sets with Colormapped Values.
Indeed, some token sets are highly individual, like the ‘anonymous’ voter who owns a set of REALTOKENS (a form of tokenized, fractional property ownership) with specific addresses in Detroit, Michigan (see Figure 16:
  1. 1.
    18481 WESTPHALIA ST DETROIT MI
  2. 2.
    9717 EVERTS ST DETROIT MI
  3. 3.
    13895 SARATOGA ST DETROIT MI
  4. 4.
    18466 FIELDING ST DETROIT MI
  5. 5.
    15095 HARTWELL ST DETROIT MI
  6. 6.
Figure 16: Map of REALTOKENS in Wallet of Anonymous Voter.
More practically, we also analyse the wallet fingerprints using Uniform Manifold Approximation and Projection (UMAP), a manifold learning technique for dimensionality reduction. UMAP seeks to reduce $D$ to a lower dimension $d$, resulting in $\mathbf{E}_{\text{reduced}}\in\mathbb{R}^{N\times d}$, while preserving the local and global structure of the data. While the UMAP algorithm is quite simple (it calculates nearest-neighbours to produce sets of weighted graphs, calculates probabilities for edges on a unified graph, and then uses a force algorithm to lay the graph out), the mathematical underpinnings are fascinating and much more complex [36]. In a sentence, UMAP finds fuzzy simplicial sets (mathematical abstractions from category theory that involve objects and mappings), which define the metric between nodes (viz. Riemannian in that space is made of individually defined metrics, unlike Euclidean space); to display a high dimensional representation in low (2 or 3) dimensions, UMAP minimizes the cross-entropy between the high-dimensional fuzzy simplicial set and the a priori low dimensional manifold.
To prepare our fingerprint dataset, we encode each voter’s unique ticker set as a sparse matrix. We then fit the wallet’s vector embeddings using a Jaccard metric with UMAP and construct an interactive 2D or 3D plot, which reveals features of the voter fingerprints (the interactive plot is available on Github). Through experimentation we found that the Jaccard, cosine, and Manhattan metrics all work reasonably well with a default 15 neighbours, but Jaccard offered the best balance of local and global topology (see Figure 18).
Figure 17: UMAP Projection of Wallet Token Fingerprints.
Figure 18: UMAP Projection of HDBSCAN Clustered Wallet Token Fingerprints.
We then cluster the embeddings using Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) and refit the embeddings with UMAP to visualize the results (see Figure 18). Inspecting the plot is revealing; for instance, when tight, uniform clusters are far from others, we suspect the presence of Sybils (clusters reveal sets of users with a highly unique wallet fingerprint; see Figure 19.
Figure 19: Suspected Sybils Determined by Wallet Fingerprints (nodes share a highly distinct fingerprint).
Finally, we train a simple deep neural network to learn each voter’s wallet token fingerprint and predict the associated DAO. While there is little practical value in this prediction, it shows that wallet sets contain important topological information. Indeed, our clustering results have an accuracy of 0.6, which is significantly better than random guessing (0.14) for seven classes, by 45.7%, although there is clearly much room for improvement.
Specifically, given a dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$ where $\mathbf{x}_{i}\in\mathbb{R}^{d}$ represents the input features of dimension $d$ and $y_{i}$ represents the labels. The dataset is divided into training, validation, and testing sets. Let $f(\mathbf{x};\theta)$ be the neural network function parameterized by $\theta$.
The neural network, $f(\mathbf{x};\theta)$, is defined as follows:
  1. 1.
    The input layer takes $d$-dimensional input vector $\mathbf{x}$.
  2. 2.
    The first linear transformation is applied: $\mathbf{z}_{1}=\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1}$, where $\mathbf{W}_{1}\in\mathbb{R}^{512\times d}$ and $\mathbf{b}_{1}\in\mathbb{R}^{512}$.
  3. 3.
    A ReLU activation function is applied: $\mathbf{a}_{1}=\text{ReLU}(\mathbf{z}_{1})$.
  4. 4.
    Batch normalization is applied to $\mathbf{a}_{1}$ before dropout: $\mathbf{a}_{1}^{\prime}=\text{BatchNorm}(\mathbf{a}_{1})$.
  5. 5.
    Dropout is applied to $\mathbf{a}_{1}^{\prime}$ with a dropout rate of 0.5.
  6. 6.
    This process is repeated for additional layers with dimensions $512\to 256\to 128\to 64$, including linear transformations, ReLU activations, batch normalization, and dropout at each step. Specifically, for layer $l$ where $2\leq l\leq 5$:
    $\mathbf{z}_{l}=\mathbf{W}_{l}\mathbf{a}_{l-1}^{\prime}+\mathbf{b}_{l},\quad\mathbf{a}_{l}^{\prime}=\text{Dropout}(\text{BatchNorm}(\text{ReLU}(\mathbf{z}_{l})),0.5)$
    with $\mathbf{W}_{l}$ and $\mathbf{b}_{l}$ representing the weights and biases of layer $l$, respectively.
  7. 7.
    The final layer’s output $\mathbf{z}_{5}$ is passed through a linear transformation without a subsequent activation or dropout for the classification: $\hat{\mathbf{y}}=\mathbf{W}_{5}\mathbf{a}_{4}^{\prime}+\mathbf{b}_{5}$, where $\mathbf{W}_{5}\in\mathbb{R}^{\text{num\_classes}\times 64}$ and $\mathbf{b}_{5}\in\mathbb{R}^{\text{num\_classes}}$.
He initialization is applied to all linear layers, ensuring variance scaling is maintained in the network’s activations. The network is trained using the cross-entropy loss function:
$\mathcal{L}(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{e^{\hat{y}_{i,y_{i}}}}{\sum_{j=1}^{\text{num\_classes}}e^{\hat{y}_{i,j}}}$
where $\hat{y}_{i,j}$ is the $j$-th element of the network’s output $\hat{\mathbf{y}}_{i}$ for input $\mathbf{x}_{i}$, and $y_{i}$ is the true label of $\mathbf{x}_{i}$.
Optimization is performed using the Adam optimizer with a learning rate of $0.001$ and weight decay of $51e-$ 5 for regularization, and the model is evaluated using a split of training, validation, and testing data.
After training, the dimensionality of embeddings is once again reduced using UMAP for inspection. We see structure, but the results are of limited value (see Figure 21). Instead, we cluster these high dimensional embeddings using a Random Forest classifier (see Figure 21).
Figure 20: UMAP Projection of Deep Feedforward Neural Network Embeddings of Wallet Token Fingerprints.
Figure 21: UMAP Projection of Random Forest Clustering of Wallet Token Fingerprints (Colored by Prediction Accuracy).
The resulting clusters can be evaluated against the true labels with a confusion matrix (Figure 22). Additionally, we computed cluster evaluation metrics (see 1).
Figure 22: Confusion Matrix of Clustering of Wallet Token Fingerprints.
Class Precision Recall F1-Score Support
0 0.52 0.28 0.37 193
1 0.60 0.60 0.60 163
2 0.46 0.38 0.42 378
3 0.44 0.38 0.41 245
4 0.60 0.67 0.63 668
5 0.47 0.36 0.41 66
6 0.61 0.72 0.66 724
Accuracy: 0.5650 (2437 samples)
Macro avg: 0.53 precision, 0.48 recall, 0.50 F1-score
Weighted avg: 0.55 precision, 0.57 recall, 0.55 F1-score
Table 1: Clustering Evaluation Metrics
While there remains much room for methodological improvement, our analysis reveals the predictive power of wallet token fingerprinting.